# Linearizability of flows by embeddings

### v6

@@ 2.2. The continuous compact case
A $C^{0}$ action of a torus $\Theta\colon T\times X\to X$ on a topological space $X$ has finitely many orbit types if there are only finitely many subgroups $H\subset T$ with the property that $H=\{g\in T\colon\Theta^{g}(x)=x\}$ for some $x\in X$ [Pal60, Def. 1.2.1].
@@
To facilitate comparison of Theorem 2 with a result of Liu, Ozay, and Sontag [LOS23a, LOS23b, Cor. 3] (see Remark 8), note that Theorem 2 remains true if “$C^{0}$ embedding” is replaced by “injective $C^{0}$ map”. This is because any injective $C^{0}$ map from a compact space to a Hausdorff space is a $C^{0}$ embedding. The same is true of Theorem 3 (by its final sentence).
@@
Theorem 2 explicitly restricts attention to $C^{0}$ torus actions with finitely many orbit types. This is not necessary in Theorem 1, since $C^{1}$ torus actions on compact manifolds always have finitely many orbit types [DK00, Prop. 2.7.1].
@@
The assumption that $X$ is homeomorphic to a subspace of a manifold has nothing to do with $\Phi$. By the Whitney embedding theorem, it is equivalent to the assumption that $X$ admits a $C^{0}$ embedding into some $\mathbb{R}^{n}$, which is clearly necessary for the existence of such a $C^{0}$ embedding that is also linearizing. The same assumption can also be reformulated intrinsically: a compact topological space $X$ is homeomorphic to a subspace of a manifold if and only if $X$ is metrizable and has finite Lebesgue covering dimension [Mun00, Thm 50.1, p. 316].
@@ 2.3. The continuous attractor basin case
Recall that a map $F\colon X\to Y$ between topological spaces is proper if $F^{-1}(K)$ is compact for every compact subset $K\subset Y$ [Lee11, p. 118].
@@
Specializing Theorem 3 to $A=\{x_{*}\}$ yields the following extension of the (globalized [LM13, EKR18, KR21]) classical Hartman-Grobman theorem [Har60a, Gro59]. The latter guarantees linearizability by a dimension-preserving $C^{0}$ embedding when $X$ is a $C^{\infty}$ manifold, $\Phi$ is the flow of a $C^{1}$ vector field, and $x_{*}$ is hyperbolic.
@@
Let $X$ be a connected and locally compact topological space homeomorphic to a subspace of a manifold. Let $\Phi$ be a $C^{0}$ flow on $X$ with a globally asymptotically stable equilibrium point $x_{*}\in X$. Then $(X,\Phi)$ is linearizable by a $C^{0}$ embedding.
Similarly, specializing Theorem 3 to the image $A$ of a limit cycle yields the following extension of the (globalized [LM13, EKR18, KR21]) classical Floquet normal form theorem ([AR67, Sec. 26], [BK94, Sec. I.3], [AMR, Sec. 4.3]), a nonlinear generalization of the classical Floquet theory of linear time-periodic systems [Hal80, Sec. III.7]. The standard Floquet normal form guarantees linearizability by a $C^{0}$ embedding that increases dimension by $1$ or $2$ when $X$ is a $C^{\infty}$ manifold, $\Phi$ is the flow of a $C^{1}$ vector field, and $x_{*}$ is hyperbolic.
@@
Let $X$ be a connected and locally compact topological space homeomorphic to a subspace of a manifold. Let $\Phi$ be a $C^{0}$ flow on $X$ with a globally asymptotically stable periodic orbit with image $A$. Then $(X,\Phi)$ is linearizable by a $C^{0}$ embedding if and only if $A$ has $C^{0}$ asymptotic phase.
@@
Similarly to Remark 5, the assumption that $X$ is locally compact and homeomorphic to a subspace of a manifold has nothing to do with $\Phi$. It is equivalent to the assumption that $X$ is homeomorphic to a locally closed subspace $S$ of some manifold $M$, meaning that $S$ is the intersection of an open subset and a closed subset of $M$ (and is equipped with the subspace topology). For example, any open or closed subset of $M$ is itself locally closed in $M$ since it is the intersection of $M$ with itself, and $M$ is both a closed and an open subset of itself. Thus, in particular, Theorem 3 applies to the restriction $\Phi$ of a flow on a manifold $M$ to the basin of attraction $X\subset M$ of an asymptotically stable compact invariant set $A\subset M$, because such a basin $X$ is always open in $M$. Finally, the same assumption can also be reformulated intrinsically: a locally compact topological space $X$ is homeomorphic to a subspace of a manifold if and only if $X$ is metrizable, separable, and has finite Lebesgue covering dimension ([Mun00, Thm 50.1, p. 316], [Eng89, Cor. 3.3.10]).
@@
Corollary 6 is complementary to a result of Liu, Ozay, and Sontag [LOS23a, LOS23b, Cor. 3]. Under some additional assumptions their result asserts that, if the omega-limit set of each $x\in X$ for $\Phi$ is nonempty and the collection of all such omega-limit sets is a countable number larger than one, then $(X,\Phi)$ is not linearizable by an injective $C^{0}$ map. Similarly, for $X$ connected Corollary 6 implies that, if $\Phi$ has more than one nonempty compact asymptotically stable invariant set, then $(X,\Phi)$ is not linearizable by an injective $C^{0}$ map. Corollary 6 and [LOS23a, LOS23b, Cor. 3] are independent since not every omega-limit set is an asymptotically stable set, and conversely.
@@ 2.4. The smooth attractor basin case
As in §2.3, if $A\subset X$ is a globally asymptotically stable compact $C^{k\geq 1}$ embedded invariant manifold for a flow $\Phi$ on a $C^{k}$ manifold $X$, we say that $A$ has $C^{k}$ asymptotic phase if a $C^{k}$ asymptotic phase map $P\colon X\to A$ for $A$ exists.
@@
    $A$ is a $C^{k}$ embedded submanifold of $X$ with $C^{k}$ asymptotic phase;
@@
Here uniquely integrable means that the vector field $x\mapsto\frac{d}{dt}\Phi^{t}(x)|_{t=0}$ has unique maximal trajectories. By the Picard-Lindelöf theorem this holds if $k\geq 2$ or, more generally, if the vector field generating $\Phi$ is $C^{1}$ or even locally Lipschitz.
@@
Let $(X,\Phi)$ be the basin of attraction of a stable hyperbolic limit cycle for a $C^{\infty}$ flow. In this case, it is known that Conditions 1, 2 are always satisfied with $k=\infty$. Thus, such a flow $(X,\Phi)$ is linearizable by a $C^{\infty}$ embedding if and only if Condition 3 is also satisfied with $k=\infty$. This is the case, e.g., under the typically-satisfied [KHR21, Thm 5] condition that the Floquet multipliers associated with the linearized flow are nonresonant [KR21, Prop. 3]. Note that, if $X$ is nonorientable, then any linearizing $C^{0}$ embedding must be dimension-increasing for this example. (Remark 12 discusses Condition 3 further.)
@@
Note that $X$ is connected if and only if $A$ is connected [Gob01, Thm 6.3]. If $X$ is not connected, then Condition 1 should be understood in the sense that the (finitely many) connected components of $A$ are compact $C^{k}$ embedded submanifolds of possibly different dimensions.
@@
If $(X,\Phi)$ is linearizable by a $C^{k\geq 1}$ embedding, then $A$ must be a normally hyperbolic invariant manifold (NHIM). More precisely, $A$ is an eventually relatively $\infty$-NHIM ([HPS77, p. 4], [EKR18, p. 4207]) with respect to any Riemannian metric on $X$. This readily follows from the proof of Theorem 4, which shows that $(X,\Phi)$ is linearizable by a proper $C^{k}$ embedding $F\colon X\to\mathbb{R}^{n}$ such that $F(A)$ is contained in the real invariant subspace for the matrix generating the linear flow corresponding to eigenspaces of purely imaginary eigenvalues. The same reasoning implies that $A$ satisfies “center bunching” conditions [EKR18, Eq. (11)] of all orders, consistent with the existence of $C^{k}$ asymptotic phase implied by such center bunching conditions [EKR18, Cor. 2].
@@
    The “stable vector bundle” ([HPS77, p. 1], [EKR18, p. 4207]) $E^{s}\subset T_{A}X$ of $A$ is $C^{k}$ and globally trivializable,
@@
    $T\Phi|_{E^{s}}$ is “$C^{k}$ reducible” (conjugate via a $C^{k}$ vector bundle automorphism covering $\textnormal{id}_{A}$) to the product of $\Phi|_{\mathbb{R}\times A}$ with a constant (with respect to some global trivialization of $E^{s}$) linear map.
Regarding (a), note that $E^{s}$ is globally trivializable if and only if $A$ is a level set of some $C^{k}$ submersion, and a sufficient condition implying $E^{s}$ is $C^{k}$ is that $A$ satisfies “$k$-center bunching” conditions [EKR18, Eq. 11]. Assuming $\Phi$ is $C^{k+1}$, sufficient conditions for (b) arise from smooth local linearization theorems for NHIMs [Ste57, Tak71, Rob71, Sel84, Sel83, Sak94, BK94] combined with a globalization technique [EKR18, Thm 2]. However, even in the case of Sternberg’s linearization theorem for a stable hyperbolic equilibrium [Ste57], the available NHIM linearization theorems yield sufficient conditions for (b) that are not necessary in general. Since Condition 2 implies that $A$ decomposes into quasiperiodic invariant tori in a certain way, sufficient conditions for (b) closest to necessary for our purposes might be obtained via techniques from the literature on normally hyperbolic invariant quasiperiodic tori [HdlL06, HCF+16] (cf. [HdlL06, Thm 4.1]). Similarly, useful sufficient conditions for (c) might be obtained via techniques from the related literature on reducibility of linear flows on vector bundles over quasiperiodic tori [JS81, Pui02] (cf. [JS81, Thm 1]). However, Condition 3 can hold even if (b), (c) do not. What is the gap between Condition 3 and (b), (c)?
@@ 3. Proofs
First assume that $\Phi$ is a $1$-parameter subgroup of a $C^{k}$ ($C^{0}$ with finitely many orbit types) action $\Theta\colon H\times X\to X$ of a torus $H$ on $X$. By the $C^{k}$ version [DK00, Thm 4.6.6.] ($C^{0}$ version [Mos57, Thm 6.1]) of the Mostow-Palais equivariant embedding theorem [Mos57, Pal57, Pal60], there exists $n\in\mathbb{N}$, a $C^{k}$ ($C^{0}$) embedding $F\colon X\to\mathbb{R}^{n}$, and a $C^{\infty}$ Lie group homomorphism $\rho\colon H\to\text{GL}(\mathbb{R}^{n})$ such that
@@
for all $t\in\mathbb{R}$, where $\rho_{*}$ is the $\rho$-induced Lie algebra homomorphism [DK00, Def. 1.10.2] and the third equality follows from $\rho$ being a group homomorphism [DK00, Lem. 1.5.1]. Thus, $F$ is a $C^{k}$ ($C^{0}$) embedding of $(X,\Phi)$ into the linear flow generated by the linear vector field $B\coloneqq\rho_{*}\boldsymbol{\omega}$.
Next, assume that $(X,\Phi)$ is linearizable by a $C^{k}$ ($C^{0}$) embedding. Since $X$ is compact, the Jordan normal form theorem implies that any linearizing $C^{0}$ embedding of $(X,\Phi)$ sends $X$ into the real invariant linear subspaces corresponding to eigenspaces of purely imaginary eigenvalues of the generating matrix $B$ of a linear flow. Thus, after embedding the real linear space into a complex linear space via complexifying and diagonalizing $B$, we may assume that $X\subset\mathbb{C}^{n}$ is an invariant subset for the linear flow generated by the ODE
@@
on $\mathbb{C}^{n}$, where $i=\sqrt{-1}$, $z=(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}$, $\boldsymbol{\omega}=(\omega_{1},\ldots,\omega_{n})\in\mathbb{R}^{n}$, and $\text{diag}(\boldsymbol{\omega})$ is the diagonal matrix with $j$-th diagonal entry $\omega_{j}$. Define a $C^{\infty}$ action $\Theta$ of $T^{n}=\mathbb{R}^{n}/\mathbb{Z}^{n}$ on $\mathbb{C}^{n}$ by $\Theta^{\tau}(z)\coloneqq\exp(2\pi i\text{diag}(\tau))z$. Note that $X$ is a union of closures of trajectories of (4) since $X$ is a closed subset of $\mathbb{C}^{n}$, with each such closure coinciding with an orbit $\Theta^{H}(z)$ of the closure $H\subset T^{n}$ of the $1$-parameter subgroup $\boldsymbol{\omega}\mathbb{R}\mod 1$. Thus, $X$ is a union of $H$-orbits, so the $C^{\infty}$ $H$ action on $\mathbb{C}^{n}$ restricts to a well-defined $C^{k}$ ($C^{0}$ with finitely many orbit types [Pal60, Thm 1.8.4]) $H$ action on $X$ with $\Phi$ a $1$-parameter subgroup of the restricted $H$ action. Finally, the closed subgroup theorem [DK00, Cor. 1.10.7] and the classification of abelian Lie groups [DK00, Cor. 1.12.4] imply that $H$ is a $C^{\infty}$ Lie subgroup of $T^{n}$ isomorphic to a torus. This completes the proof. ∎
@@
By Theorem 1, $\Phi$ is a $1$-parameter subgroup of a $C^{1}$ torus action on $X$. By restricting the torus action to the action of the closure $H$ of this $1$-parameter subgroup, we obtain a $C^{1}$ action $\Theta$ of a subtorus $H$ on $X$ with the property that the equilibria of $\Phi$ coincide with the fixed points of $\Theta$. Let $x\in X$ be an isolated such equilibrium/fixed point. By Bochner’s linearization theorem, there is a $\Theta$-invariant open neigborhood $U$ of $x$ and $C^{1}$ local coordinates on $U$ in which $\Theta$ is the linear action of a closed subgroup of the orthogonal group $O(n)$ on $\mathbb{R}^{n}$ ([DK00, Thm 2.2.1], [Hir94, Thm 4.7.1]). Thus, in these coordinates $\Phi$ corresponds to the action of a $1$-parameter subgroup of $O(n)$, and hence the vector field generating $\Phi$ corresponds to a skew-symmetric linear vector field $B$. Moreover, $B$ is invertible since $x$ is an isolated equilibrium of $\Phi$ and hence an isolated zero of $B$. The fact that all invertible skew-symmetric matrices have even dimensions implies that $\dim X$ is even. Finally, the determinant of $B$ is positive since eigenvalues of an invertible real skew-symmetric matrix come in imaginary conjugate pairs, so the Hopf index of the vector field generating $\Phi$ at $x$ is equal to $1$. ∎
@@
The condition $F\circ\Phi^{t}=\boldsymbol{\omega}t+F\mod 1$ implies that the open set $R\subset T^{n}$ of regular values of $F$ is invariant under the irrational linear flow $(t,y)\mapsto\boldsymbol{\omega}t+y\mod 1$. Since an irrational linear flow is minimal, either $R=T^{n}$ or $R=\varnothing$. Sard’s theorem [Hir94, Thm 3.1.3] eliminates the latter option, so $F$ is a local diffeomorphism, hence also a covering map since $X$ is compact [Lee11, p. 303, 11-9]. The classification of covering spaces of $T^{n}$ [Hat02, §1.3] yields a natural $C^{k}$ identification of $X$ with $T^{n}=\mathbb{R}^{n}/\mathbb{Z}^{n}$ such that $F\colon\mathbb{R}^{n}/\mathbb{Z}^{n}\approx X\to\mathbb{R}^{n}/\mathbb{Z}^{n}$ is additionally a Lie group homomorphism, so is given by an invertible matrix $B$ with integer entries. Since $F\circ\Phi^{t}=\boldsymbol{\omega}t+F\mod 1$, this implies that $\Phi$ is identified with the flow of the constant vector field $\tilde{\boldsymbol{\omega}}=B^{-1}\boldsymbol{\omega}$. The entries of $B^{-1}$ are rational, so rational independence of the components of $\boldsymbol{\omega}$ implies the same for $\tilde{\boldsymbol{\omega}}$. ∎
@@
Since $X$ is locally compact and $A$ is globally asymptotically stable, there exists a proper $C^{0}$ Lyapunov function $V\colon X\to[0,\infty)$ satisfying $V^{-1}(0)=A$ and strictly decreasing along $\Phi$-trajectories outside $A$ [BS67, Thm 2.7.20]. Fix any $c\in(0,\infty)$ and define the compact set $N\coloneqq V^{-1}(c)$. Each trajectory in $X\setminus A$ crosses $N$ exactly once, and the “time-to-impact-N” map $\tau\colon X\setminus A\to(-\infty,\infty)$ is $C^{0}$ (by the method of proof of [BS67, Thm 2.7.14] or [Con78, Theorem II.2.3]).

### v7

@@ 2.2. The continuous compact case
A $C^{0}$ action of a torus $\Theta\colon T\times X\to X$ on a topological space $X$ has finitely many orbit types if there are only finitely many subgroups $H\subset T$ with the property that $H=\{g\in T\colon\Theta^{g}(x)=x\}$ for some $x\in X$ [52, Def. 1.2.1].
@@
To facilitate comparison of Theorem 2 with a result of Liu, Ozay, and Sontag [40, 41, Cor. 3] (see Remark 9), note that Theorem 2 remains true if $C^{0}$ embedding” is replaced by “injective $C^{0}$ map”. (This is because any injective $C^{0}$ map from a compact space to a Hausdorff space is a $C^{0}$ embedding.) The same is true of Theorem 3 (by its final sentence).
@@
Theorem 2 explicitly restricts attention to $C^{0}$ torus actions with finitely many orbit types. This is not necessary in Theorem 1, since $C^{1}$ torus actions on compact manifolds always have finitely many orbit types [53, Prop. 2.7.1].
@@
The assumption that $X$ is homeomorphic to a subspace of a manifold has nothing to do with $\Phi$. By the Whitney embedding theorem, it is equivalent to the assumption that $X$ admits a $C^{0}$ embedding into some $\mathbb{R}^{n}$, which is clearly necessary for the existence of such a $C^{0}$ embedding that is also linearizing. The same assumption can also be reformulated intrinsically: a compact topological space $X$ is homeomorphic to a subspace of a manifold if and only if $X$ is metrizable and has finite Lebesgue covering dimension [54, Thm 50.1, p. 316].
@@ 2.3. The continuous attractor basin case
Recall that a map $F\colon X\to Y$ between topological spaces is proper if $F^{-1}(K)$ is compact for every compact subset $K\subset Y$ [1, p. 118].
@@
Specializing Theorem 3 to $A=\{x_{*}\}$ yields the following extension of the (globalized [27, 28, 29, 55]) classical Hartman-Grobman theorem [18, 19]. The latter guarantees linearizability by a dimension-preserving $C^{0}$ embedding when $X$ is a $C^{\infty}$ manifold, $\Phi$ is the flow of a $C^{1}$ vector field, and $x_{*}$ is hyperbolic.
@@
Let $X$ be a locally compact topological space homeomorphic to a subspace of a manifold. Let $\Phi$ be a $C^{0}$ flow on $X$ with a globally asymptotically stable equilibrium point $x_{*}\in X$. Then $(X,\Phi)$ is linearizable by a $C^{0}$ embedding.
Similarly, specializing Theorem 3 to the image $A$ of a limit cycle yields the following extension of the (globalized [27, 28, 29, 55]) classical Floquet normal form theorem ([56, Sec. 26], [44, Sec. I.3], [57, Sec. 4.3]), a nonlinear generalization of the classical Floquet theory of linear time-periodic systems [45, Sec. III.7]. The standard Floquet normal form guarantees linearizability by a $C^{0}$ embedding when $X$ is a $C^{\infty}$ manifold, $\Phi$ is the flow of a $C^{1}$ vector field, and $A$ is normally hyperbolic.
@@
Let $X$ be a locally compact topological space homeomorphic to a subspace of a manifold. Let $\Phi$ be a $C^{0}$ flow on $X$ with a globally asymptotically stable periodic orbit with image $A$. Then $(X,\Phi)$ is linearizable by a $C^{0}$ embedding if and only if $A$ has $C^{0}$ asymptotic phase.
@@
Similarly to Remark 6, the assumption that $X$ is locally compact and homeomorphic to a subspace of a manifold has nothing to do with $\Phi$. It is equivalent to the assumption that $X$ is homeomorphic to a locally closed subspace $S$ of some manifold $M$, meaning that $S$ is the intersection of an open subset and a closed subset of $M$ (and is equipped with the subspace topology). For example, any open or closed subset of $M$ is itself locally closed in $M$ since it is the intersection of $M$ with itself, and $M$ is both a closed and an open subset of itself. Thus, in particular, Theorem 3 applies to the restriction $\Phi$ of a flow on a manifold $M$ to the basin of attraction $X\subset M$ of an asymptotically stable compact invariant set $A\subset M$, because such a basin $X$ is always open in $M$. Finally, the same assumption can also be reformulated intrinsically: a locally compact topological space $X$ is homeomorphic to a subspace of a manifold if and only if $X$ is metrizable, separable, and has finite Lebesgue covering dimension ([54, Thm 50.1, p. 316], [58, Cor. 3.3.10]).
@@
Corollary 6 is complementary to a result of Liu, Ozay, and Sontag [40, 41, Cor. 3]. Under some additional assumptions their result asserts that, if all forward $\Phi$-orbits are precompact and the collection of all omega-limit sets of points $x\in X$ is countable and contains more than one element, then $(X,\Phi)$ is not linearizable by an injective $C^{0}$ map. Similarly, for $X$ connected Corollary 6 implies that, if $\Phi$ has more than one nonempty compact asymptotically stable invariant set, then $(X,\Phi)$ is not linearizable by an injective $C^{0}$ map. Corollary 6 and [40, 41, Cor. 3] are independent since not every omega-limit set is an asymptotically stable set, and conversely.
@@ 2.4. The smooth attractor basin case
As in §2.3, if $A\subset X$ is a globally asymptotically stable compact $C^{k}$ embedded invariant manifold for a flow $\Phi$ on a $C^{k}$ manifold $X$, we say that $A$ has $C^{k}$ asymptotic phase if a $C^{k}$ asymptotic phase map $P\colon X\to A$ for $A$ exists.
@@
    AA is a $C^{k}$ embedded submanifold of $X$ with $C^{k}$ asymptotic phase;
@@
Here uniquely integrable means that the vector field $x\mapsto\frac{d}{dt}\Phi^{t}(x)|_{t=0}$ has unique maximal trajectories. By the Picard-Lindelöf theorem this holds if $k\geq 2$ or, more generally, if the vector field generating $\Phi$ is locally Lipschitz.
@@
Let $(X,\Phi)$ be the basin of attraction of a stable hyperbolic limit cycle for a $C^{\infty}$ flow. In this case, it is known that Conditions 1, 2 are always satisfied with $k=\infty$. Thus, such a flow $(X,\Phi)$ is linearizable by a $C^{\infty}$ embedding if and only if Condition 3 is also satisfied with $k=\infty$. This is the case, e.g., under the typically-satisfied condition that the Floquet multipliers associated with the linearized flow are nonresonant [55, Prop. 3]. Note that, if $X$ is nonorientable, then any linearizing $C^{0}$ embedding must be dimension-increasing for this example. (Remark 13 discusses Condition 3 further.)
@@
Note that $X$ is connected if and only if $A$ is connected [59, Thm 6.3]. If $X$ is not connected, then Condition 1 should be understood in the sense that the (finitely many) connected components of $A$ are compact $C^{k}$ embedded submanifolds of possibly different dimensions.
@@
If $(X,\Phi)$ is linearizable by a $C^{k\geq 1}$ embedding, then $A$ must be a normally hyperbolic invariant manifold (NHIM). More precisely, $A$ is an eventually relatively $\infty$-NHIM ([43, p. 4], [29, p. 4207]) with respect to any Riemannian metric on $X$. This follows readily from the proof of Theorem 4, which shows that $(X,\Phi)$ is linearizable by a proper $C^{k}$ embedding $F\colon X\to\mathbb{R}^{n}$ such that $F(A)$ is contained in the real invariant subspace for the matrix generating the linear flow corresponding to eigenspaces of purely imaginary eigenvalues. The same reasoning implies that $A$ satisfies “center bunching” conditions [29, Eq. (11)] of all orders, consistent with the existence of $C^{k}$ asymptotic phase implied by such center bunching conditions [29, Cor. 2].
@@
    The “stable vector bundle” ([43, p. 1], [29, p. 4207]) $E^{s}\subset T_{A}X$ of $A$ is $C^{k}$ and globally trivializable,
@@
    $T\Phi|_{E^{s}}$ is $C^{k}$ reducible” (conjugate via a $C^{k}$ vector bundle automorphism covering $\textnormal{id}_{A}$) to the product of $\Phi|_{\mathbb{R}\times A}$ with a constant (with respect to some global trivialization of $E^{s}$) linear map.
Regarding (a), note that $E^{s}$ is globally trivializable if and only if $A$ is a level set of some $C^{k}$ submersion, and a sufficient condition implying $E^{s}$ is $C^{k}$ is that $A$ satisfies “$k$-center bunching” conditions [29, Eq. 11]. Assuming $\Phi$ is $C^{k+1}$, sufficient conditions for (b) arise from $C^{k}$ local linearization theorems for NHIMs [16, 23, 22, 25, 24, 26, 44] combined with a globalization technique [28, sec. 4.6] (see also [29, Thm 2]). However, even in the case of Sternberg’s linearization theorem for a stable hyperbolic equilibrium [16], the available NHIM linearization theorems yield sufficient conditions for (b) that are not necessary in general. Since Condition 2 implies that $A$ decomposes into quasiperiodic invariant tori in a certain way, sufficient conditions for (b) closest to necessary for our purposes might be obtained via techniques from the literature on normally hyperbolic invariant quasiperiodic tori [60, 61] (cf. [60, Thm 4.1], [5, Sec. 9]). Similarly, useful sufficient conditions for (c) might be obtained via techniques from the related literature on reducibility of linear flows on vector bundles over quasiperiodic tori [62, 63] (cf. [62, Thm 1]). However, Condition 3 can hold even if (b), (c) do not. What is the gap between Condition 3 and (b), (c)?
@@ 3. Proofs
First assume that $\Phi$ is a $1$-parameter subgroup of a $C^{k}$ ($C^{0}$ with finitely many orbit types) action $\Theta\colon H\times X\to X$ of a torus $H$ on $X$. By the $C^{k}$ version [53, Thm 4.6.6.] ($C^{0}$ version [64, Thm 6.1]) of the Mostow-Palais equivariant embedding theorem [64, 65, 52], there exists $n\in\mathbb{N}$, a $C^{k}$ ($C^{0}$) embedding $F\colon X\to\mathbb{R}^{n}$, and a $C^{\infty}$ Lie group homomorphism $\rho\colon H\to\text{GL}(\mathbb{R}^{n})$ such that
@@
for all $t\in\mathbb{R}$, where $\rho_{*}$ is the $\rho$-induced Lie algebra homomorphism [53, Def. 1.10.2] and the third equality follows from $\rho$ being a group homomorphism [53, Lem. 1.5.1]. Thus, $F$ is a $C^{k}$ ($C^{0}$) embedding of $(X,\Phi)$ into the linear flow generated by the linear vector field $B\coloneqq\rho_{*}\boldsymbol{\omega}$.
Next, assume that $(X,\Phi)$ is linearizable by a $C^{k}$ ($C^{0}$) embedding. Since $X$ is compact, the Jordan normal form theorem implies that any linearizing $C^{0}$ embedding of $(X,\Phi)$ sends $X$ into the sum of real invariant linear subspaces corresponding to eigenspaces of purely imaginary eigenvalues of the matrix $B$ generating the linear flow. Thus, after embedding the real linear space into a complex linear space via complexifying and diagonalizing $B$, we may assume that $X\subset\mathbb{C}^{n}$ is an invariant subset for the linear flow generated by the ODE
@@
on $\mathbb{C}^{n}$, where $i=\sqrt{-1}$, $z=(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}$, $\boldsymbol{\omega}=(\omega_{1},\ldots,\omega_{n})\in\mathbb{R}^{n}$, and $\text{diag}(\boldsymbol{\omega})$ is the diagonal matrix with $j$-th diagonal entry $\omega_{j}$. Define a $C^{\infty}$ action $\Theta$ of $T^{n}=\mathbb{R}^{n}/\mathbb{Z}^{n}$ on $\mathbb{C}^{n}$ by $\Theta^{\tau}(z)\coloneqq\exp(2\pi i\text{diag}(\tau))z$. Note that $X$ is a union of closures of trajectories of (4) since $X$ is a closed subset of $\mathbb{C}^{n}$, with each such closure coinciding with an orbit $\Theta^{H}(z)$ of the closure $H\subset T^{n}$ of the $1$-parameter subgroup $\boldsymbol{\omega}\mathbb{R}\mod 1$. Thus, $X$ is a union of $H$-orbits, so the $C^{\infty}$ $H$ action on $\mathbb{C}^{n}$ restricts to a well-defined $C^{k}$ ($C^{0}$ with finitely many orbit types [52, Thm 1.8.4]) $H$ action on $X$ with $\Phi$ a $1$-parameter subgroup of the restricted $H$ action. Finally, the closed subgroup theorem [53, Cor. 1.10.7] and the classification of abelian Lie groups [53, Cor. 1.12.4] imply that $H$ is a $C^{\infty}$ Lie subgroup of $T^{n}$ isomorphic to a torus. This completes the proof. ∎
@@
By Theorem 1, $\Phi$ is a $1$-parameter subgroup of a $C^{1}$ torus action on $X$. By restricting the torus action to the action of the closure $H$ of this $1$-parameter subgroup, we obtain a $C^{1}$ action $\Theta$ of a subtorus $H$ on $X$ with the property that the equilibria of $\Phi$ coincide with the fixed points of $\Theta$. Let $x\in X$ be an isolated such equilibrium/fixed point. By Bochner’s linearization theorem, there is a $\Theta$-invariant open neighborhood $U$ of $x$ and $C^{1}$ local coordinates on $U$ in which $\Theta$ is the linear action of a closed subgroup of the orthogonal group $O(n)$ on $\mathbb{R}^{n}$ ([53, Thm 2.2.1], [2, Thm 4.7.1]). Thus, in these coordinates $\Phi$ corresponds to the action of a $1$-parameter subgroup of $O(n)$, and hence the vector field generating $\Phi$ corresponds to a skew-symmetric linear vector field $B$. Moreover, $B$ is invertible since $x$ is an isolated equilibrium of $\Phi$ and hence an isolated zero of $B$. The fact that all invertible skew-symmetric matrices have even dimensions implies that $\dim X$ is even. Finally, the determinant of $B$ is positive since eigenvalues of an invertible real skew-symmetric matrix come in imaginary conjugate pairs, so the Hopf index of the vector field generating $\Phi$ at $x$ is equal to $1$. ∎
@@
The condition $F\circ\Phi^{t}=\boldsymbol{\omega}t+F\mod 1$ implies that the open set $R\subset T^{n}$ of regular values of $F$ is invariant under the irrational linear flow $(t,y)\mapsto\boldsymbol{\omega}t+y\mod 1$. Since an irrational linear flow is minimal, either $R=T^{n}$ or $R=\varnothing$. Sard’s theorem [2, Thm 3.1.3] eliminates the latter option, so $F$ is a local diffeomorphism, hence also a covering map since $X$ is compact [1, p. 303, 11-9]. The classification of covering spaces of $T^{n}$ [66, §1.3] yields a natural $C^{k}$ identification of $X$ with $T^{n}=\mathbb{R}^{n}/\mathbb{Z}^{n}$ such that $F\colon\mathbb{R}^{n}/\mathbb{Z}^{n}\approx X\to\mathbb{R}^{n}/\mathbb{Z}^{n}$ is additionally a Lie group homomorphism, so is given by an invertible matrix $B$ with integer entries. Since $F\circ\Phi^{t}=\boldsymbol{\omega}t+F\mod 1$, this implies that $\Phi$ is identified with the flow of the constant vector field $\tilde{\boldsymbol{\omega}}=B^{-1}\boldsymbol{\omega}$. The entries of $B^{-1}$ are rational, so rational independence of the components of $\boldsymbol{\omega}$ implies the same for $\tilde{\boldsymbol{\omega}}$. ∎
@@
Since $X$ is locally compact and $A$ is globally asymptotically stable, there exists a proper $C^{0}$ Lyapunov function $V\colon X\to[0,\infty)$ satisfying $V^{-1}(0)=A$ and strictly decreasing along $\Phi$-trajectories outside $A$ [67, Thm 2.7.20]. Fix any $c\in(0,\infty)$ and define the compact set $N\coloneqq V^{-1}(c)$. Each trajectory in $X\setminus A$ crosses $N$ exactly once, and the “time-to-impact-NN” map $\tau\colon X\setminus A\to(-\infty,\infty)$ is $C^{0}$ (by the method of proof of [67, Thm 2.7.14] or [68, Theorem II.2.3]).